OpenStudy (unofficialllyy):

On a distant planet a freely falling object has an acceleration of 25 m/s2 . Calculate the vertical distance an object dropped from rest on this planet covers in 1.5 s .

OpenStudy (raffle_snaffle):

delta_y = vi*t + 1/2at^2

OpenStudy (raffle_snaffle):

plug and play.

OpenStudy (unofficialllyy):

What is delta_y?

OpenStudy (raffle_snaffle):

the change in the height

TheSmartOne (thesmartone):

\[\lim_{t \rightarrow 1.5}\frac{ dx }{ dt }\] where x = 1/2at^2 first find the derivative and then plug in t = 1.5 and remember a = 25 m/s^2 since we're on a distant planet and not Earth

OpenStudy (unofficialllyy):

I got 28 m

OpenStudy (raffle_snaffle):

yep.

TheSmartOne (thesmartone):

That's wrong.